OpenStudy (anonymous):

cscθ-cotθ=sinθ/(1+cosθ) im working on the left side and i got: (1/sinθ) - cosθ/sinθ = (1- cosθ)/sinθ now im stuck. help?

OpenStudy (anonymous):

what did you do to the right? What is the assignment, are you proving one side equal to other, or solving equation?

OpenStudy (rob1525):

what the original problem?

OpenStudy (anonymous):

yep, that works... keep going:)

OpenStudy (anonymous):

i am proving that both sides are equal and im only allowed to work on one side.

OpenStudy (anonymous):

\[\frac{ 1-\cos \theta }{\sin \theta } =\frac {\sin \theta } {1+\cos \theta }\]

OpenStudy (anonymous):

oh, only allowed to work on the LHS?

OpenStudy (anonymous):

that is what i have so far but im stuck and what does LHS mean?

OpenStudy (anonymous):

left hand side

OpenStudy (anonymous):

i can work on either side but once i start working on one side im limited to that side.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

use sin^2 x + cos ^2 x =1

OpenStudy (anonymous):

sin^2 x = 1- cos^2 x = (1+cosx)(1-cosx)

OpenStudy (rob1525):

multiply the top and bottom by cos(theta) .right, then use identities to simplify.

OpenStudy (anonymous):

sin x / (1-cosx) = (1+cosx)/sin x

OpenStudy (anonymous):

or vice versa:)

OpenStudy (anonymous):

It requires algebraic manipulation. Continuing from Algebraic work above